what is the probability of getting either a nine or a queen, when drawing a single card from a deck of 52 cards?

Answer by checkley79(3341)

Probability=success/total events.
There are 4 nines & 4 queens.
P=(4+4)/52
p=8/52=.1538 or 15.38% ans.